OpenStudy (anonymous):

An electric company charges $0.018 per kilowatt hour (kWh) for the first 500 kWh, $0.056 for the kWhs over 500 kWh. Write a piece wise function that best represents the amount of a bill based on the total kWh, x ?

OpenStudy (anonymous):

thx guys

OpenStudy (ahvei):

huh?

OpenStudy (bibby):

\(f(x) = \begin{cases}? & 0 \le x \le 500 \\ ? & \text{otherwise} \end{cases} \)

OpenStudy (anonymous):

is that all

OpenStudy (anonymous):

this is haleys hw

OpenStudy (anonymous):

your on flvs

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we were talking about ur toes

OpenStudy (anonymous):

help her with the question marks

OpenStudy (bibby):

the way it works is you're just multiplying the rate per hour times the amount they used so if it's under 500kWh, you pay 0.018x if above 500 you pay 0.056x

OpenStudy (anonymous):

wotm8

OpenStudy (anonymous):

so all yuo do is multiply the rate per hour by the amount they used

OpenStudy (anonymous):

is the numbers hat the question marks are

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

i will jus help

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

is that the answe

OpenStudy (anonymous):

tank u com agan